You are a network administrator and you have the following address 172.31.0.0/16. You have to logically divide it to accommodate the following:

Subnet 1: 5,000 PCs, 100 printers, 10 servers and 21 IP phones

Subnet 2: 2,500 PCs, 45 printers, 4 servers and 100 IP phones

Subnet 3: 1,500 PCs, 75 printers, 200 scanners, 25 servers and 250 IP phones

Subnet 4: 1,000 PCs, 250 servers, 250 IP phones

Subnet 5: 750 PCs, 50 servers, 50 IP phones and 25 scanners

How many bits will you borrow?

How many subnets and hosts will there be?

List the usable subnets with the new subnet mask.

Does the addressing scheme have the ability to grow as the network grows?
Answer #1

Solution for the problem is provided below, please comment if any doubts:

Total number of hosts are there in each subnets are:

Subnet 1: 5000 + 100 + 10 + 21 = 5131

Subnet 2: 2500 + 45 + 4 + 100 = 2649

Subnet 3: 1500 + 75 + 200 + 25 + 250 = 2050

Subnet 4: 1000 + 250 + 250 = 1500

Subnet 5: 750 + 50 + 50 + 25 = 875

Number of bits required to address the highest number host subnet = Ceil(log2(5131)) = 13

Thus remaining bits that are available in the total available bits allotted for the hosts = 16 – 13 = 3

Thus number of bits borrow = 3

Number bits that can be used to address the subnets = 3

Total number of subnets = 23 = 8

Each subnets have 13 bits to address the hosts, thus total number of hosts = 8*213 = 65536

Total number of hosts = 65536

Subnet

Subnet mask

172.31.0.0/13

255.255.224.0

172.31.32.0/13

255.255.224.0

172.31.64.0/13

255.255.224.0

172.31.96.0/13

255.255.224.0

172.31.128.0/13

255.255.224.0

172.31.160.0/13

255.255.224.0

172.31. 192.0/13

255.255.224.0

172.31.224.0/13

255.255.224.0

Yes the addressing schema has the ability to grow as the network grows, since:

  • In the allotted 8 subnets only five are used, rest three is unused, thus it can be used to allot the network when it grows. Also there are unused host addresses for each subnet used also, thus it is scalable in that respect also.